Place the correct symbol ( <, >, = ) for the following expressions.
|34| __ |-34|
|-10| __ |5|
|-2/3| __ |2/3|
|-34| __ -|-34|

Answer by tutorcecilia(2152)

|34| = |-34| because |34| = 34 and the |-34|also = 34
|-10| > |5| because |-10| = 10 and the |5| = 5 and 10>5
|-2/3| = |2/3| because |-2/3| = 2/3 and |2/3| also =2/3
|-34| > -|-34|because |-34| = 34 and -|-34|= -(34)=-34
In the last problem, the absolute value was multiplied by (-1).